### v1

Abstract
This paper investigates a Halpern acceleration of the inexact proximal point method for solving maximal monotone inclusion problems in Hilbert spaces. The proposed Halpern inexact proximal point method (HiPPM) is shown to be globally convergent, and a unified framework is developed to analyze its worst-case convergence rate. Under mild summability conditions on the inexactness tolerances, HiPPM achieves an $\mathcal{O}(1/k^{2})$ rate in terms of the squared fixed-point residual. Furthermore, under additional mild condition, the method retains a fast linear convergence rate. Building upon this framework, we further extend the acceleration technique to constrained convex optimization through the augmented Lagrangian formulation. In analogy to Rockafellar’s classical results, the resulting accelerated inexact augmented Lagrangian method inherits the convergence rate and complexity guarantees of HiPPM. The analysis thus provides a unified theoretical foundation for accelerated inexact proximal algorithms and their augmented Lagrangian extensions.
Key words: proximal point method, Halpern acceleration, inexactness, global convergence, rate of convergence.
AMS subject classification: 90C30
1 Introduction
[…]
$\text{find }z\in\mathcal{H}\text{ such that }0\in T(z),$ (1.1)
where $T:\mathcal{H}\rightrightarrows\mathcal{H}$ is a maximally monotone operator. This problem provides a unifying framework for a broad range of models in optimization and equilibrium theory. In particular, many convex optimization problems, variational inequalities, and saddle-point formulations can be recast as finding the zeros of a monotone operator. The proximal point method (PPM) is a fundamental approach for solving monotone inclusion problems of the form (1.1). It not only enjoys strong global convergence guarantees for maximal monotone operators, but also serves as the theoretical foundation for various operator-splitting schemes. In particular, for convex optimization problems, both the augmented Lagrangian method (ALM) and the alternating direction method of multipliers (ADMM) can be interpreted as instances of the PPA when applied to their respective dual formulations.
The proximal mapping associated with parameter $c>0$, denoted by $P_{c}$ is defined by
$P_{c}(z)=(I+cT)^{-1}(z),z\in{\cal H}.$ (1.2)
It follows from [10, Proposition 1] that $P_{c}$ is a nonexpansive mapping with
$0\in T(z)\mbox{ if and only if }z=P_{c}(z),$
then or equivalently,
$0\in T(z)\mbox{ if and only if }Q_{c}(z)=0,$
where $Q_{c}=I-P_{c}$. The classical proximal point method is based on solving a fixed point of $P_{c}$, which can be stated as follows
[…]
The inexact proximal point method of Rockafellar [10] is of the following form
[…]
Two rules for the approximate calculation of $P_{c_{k}}(z^{k})$ are criteria (A) and (B), described as follows
$\begin{array}[]{l}(A)\quad\quad\quad\quad\quad\quad\|z^{k+1}-P_{c_{k}}(z^{k})\|\leq\varepsilon_{k},\quad\displaystyle\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty;\\[6.0pt] (B)\quad\quad\quad\quad\quad\quad\|z^{k+1}-P_{c_{k}}(z^{k})\|\leq\delta_{k}\|z^{k+1}-z^{k}\|,\quad\displaystyle\sum_{k=0}^{\infty}\delta_{k}<+\infty.\end{array}$ (1.5)
It was shown in [1] that, for maximum monotone operator inclusion problems without any regularity assumptions, the proximal point method admits a worst-case convergence rate of $\mathcal{O}(1/k)$, measured in terms of the squared fixed-point residual. A tighter sublinear convergence rate was later established as $\mathcal{O}\left(1/k(1+(k-1)^{-1})^{k-1}\right)$ in [3] when the underlying Euclidean space has dimension greater than or equal to two. To further enhance the efficiency of the proximal point method, accelerating its worst-case performance has become a significant focus in both theoretical research and practical applications. In particular, Nesterov-type acceleration [7, 8] has been shown to improve the rate to $\mathcal{O}(1/k^{2})$. For example, Xu [12] proposed an inexact accelerated augmented Lagrangian method for constrained convex programming.
In contrast to Nesterov-type acceleration, which is originally developed for convex optimization, Halpern’s fixed-point iteration [4] is designed for approximating fixed points of nonexpansive operators. Since Contreras and Cominetti [2] revealed a close connection between Kim’s accelerated proximal point method [5] and Halpern’s iteration, it is natural to consider Halpern-type schemes for accelerating the proximal point method, given their simplicity. Lieder [6] studied Halpern’s iteration in Hilbert spaces and establish the $\mathcal{O}(1/k^{2})$ convergence rate. Moreover, under the co-coercive assumption, Tran-Dinh [11] derive an equivalence between the Nesterov’s accelerate method and Halpern’s fixed-point iteration. However, these connections do not directly extend to settings in which the subproblems are solved inexactly, which is typically the case in practice.
In this paper, we study a Halpern-type acceleration of the inexact proximal point method (PPM) for solving the maximal monotone inclusion problem (1.1) in Hilbert spaces. The main contributions of this work can be summarized as follows:
  • 1.
    We establish the global convergence of the proposed inexact PPM with Halpern iteration and provide a unified framework for analyzing its convergence rate. Under mild conditions on the inexactness tolerance sequence, the squared fixed-point residual is shown to converge at the rate of $\mathcal{O}(1/k^{2})$.
  • 2.
    As pointed out in [5], for strongly monotone operators, the standard proximal point method enjoys a linear convergence rate in terms of the fixed-point residual, whereas its Halpern-accelerated variant is not necessarily guaranteed to preserve such a rate. In this paper, we demonstrate that the inexact PPM with Halpern acceleration indeed attains linear convergence under the strong monotonicity assumption.
  • 3.
    Building upon the accelerated inexact proximal point framework, we further extend the analysis to constrained convex optimization problems within the augmented Lagrangian setting. In analogy with Rockafellar’s seminal work [10, 9], the proposed accelerated inexact PPM naturally induces an accelerated inexact augmented Lagrangian method. This connection allows us to derive corresponding convergence rate and iteration complexity results for the accelerated iALM.
The remainder of this paper is organized as follows. Section 2 introduces the framework of the Halpern accelerated inexact proximal point algorithm ( in Hilbert spaces and establishes its global convergence. Section 3 develops a general framework for analyzing the worst-case convergence rate of the proposed HiPPA using a direct algebraic proof, and demonstrates its fast linear convergence under mild conditions. Section 4 further derives the convergence rate of the Halpern accelerated inexact augmented Lagrangian method, based on its close connection with the Halpern accelerated inexact PPA.
Notation. Throughout the paper, $\mathcal{H}$ denotes a real Hilbert space equipped with inner product $\langle\cdot,\cdot\rangle$ and the induced norm $\|\cdot\|$. A set-valued mapping $T:\mathcal{H}\rightrightarrows\mathcal{H}$ is said to be monotone if $\langle z-z^{\prime},\,w-w^{\prime}\rangle\geq 0$ whenever $w\in T(z),\,w^{\prime}\in T(z^{\prime})$. It is called maximal monotone if, in addition, its graph ${\rm gph}T:=\{(z,w)\in\mathcal{H}\times\mathcal{H}:w\in T(z)\}$ is not properly contained in the graph of any other monotone operator $T^{\prime}:\mathcal{H}\rightrightarrows\mathcal{H}$. We denote by $\mathbb{I}_{C}$ the indicator function of a set $C$, and by ${\rm Dom}(f)$ the effective domain of a function $f$. All other notation will be defined as needed in the subsequent sections.
2 Halpern acceleration of the inexact PPA
Building upon Rockafellar’s inexact proximal point algorithm [9], we introduce an additional Halpern-type acceleration step to enhance its convergence behavior. The resulting accelerated scheme, referred to as HiPPA, can be formulated as follows.
[…]
Here criteria (A) and (B) are of the following forms:
$\begin{array}[]{l}(A)\quad\quad\quad\quad\quad\quad\|\bar{z}^{k}-P_{c_{k}}(z^{k})\|\leq\varepsilon_{k},\quad\displaystyle\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty;\\[6.0pt] (B)\quad\quad\quad\quad\quad\quad\|\bar{z}^{k}-P_{c_{k}}(z^{k})\|\leq\delta_{k}\|\bar{z}^{k}-z^{k}\|,\quad\displaystyle\sum_{k=0}^{\infty}\delta_{k}<+\infty.\end{array}$ (2.2)
Assumption 2.1.
Let $T:{\cal H}\rightrightarrows{\cal H}$ be a maximal monotone set-valued mapping and $T^{-1}(0)\neq\emptyset$, where $T^{-1}(0)=\{z\in{\cal H}:0\in T(z)\}$ is the solution set of the inclusion problem (1.1).
We recall the following result from [10, Proposition 2], which serves as a fundamental criterion for the existence of solutions to the inclusion problem $0\in T(z)$.
Lemma 2.1.
Suppose that for some $\tilde{z}\in{\cal H}$ and $\rho>0$ one has
$\langle z-\tilde{z},w\rangle\geq 0\mbox{ f or all }z,w\mbox{ with }w\in T(z),\|z-\tilde{z}\|\geq\rho.$
Then there exists at least one $z$ satisfying $0\in T(z)$. Moreover, this condition is both necessary and sufficient for the existence of a solution.
For notational convenience,, for $k=0,1,\ldots$, define
$\eta^{k}=\bar{z}^{k}-P_{c_{k}}(z^{k}).$ (2.3)
Then, for the sequence $\{z^{k}\}$ generated by Algorithm (2.1), one has that $\|\eta^{k}\|\leq\varepsilon_{k}$ when criterion (A) is adopted, and $\|\eta^{k}\|\leq\delta_{k}\|\bar{z}^{k}-z^{k}\|$ when criterion (B) is adopted.
Lemma 2.2.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is generated by Algorithm (2.1) with criterion (A). Then for any $z^{*}\in T^{-1}(0)$,
$\|z^{k}-z^{*}\|\leq\|z^{0}-z^{*}\|+\displaystyle\sum_{{j=0}}^{k-1}\varepsilon_{j},\,\,k=1,2,\ldots.$ (2.4)
Proof.
The proof proceeds by induction. For $k=0$,
$z^{1}=\displaystyle\frac{1}{2}z^{0}+\displaystyle\frac{1}{2}(P_{c_{0}}(z^{0})+\eta^{0}),$
which implies
$\begin{array}[]{ll}\|z^{1}-z^{*}\|&=\left\|\displaystyle\frac{1}{2}[z^{0}-z^{*}]+\displaystyle\frac{1}{2}[P_{c_{0}}(z^{0})-z^{*}]+\displaystyle\frac{1}{2}\eta^{0}\right\|\\[8.0pt] &\leq\displaystyle\frac{1}{2}\|z^{0}-z^{*}\|+\displaystyle\frac{1}{2}\|P_{c_{0}}(z^{0})-z^{*}\|+\displaystyle\frac{1}{2}\|\eta^{0}\|\\[8.0pt] &\leq\|z^{0}-z^{*}\|+\varepsilon_{0}.\end{array}$
Suppose the result is valid until $k\geq 1$, then
$\begin{array}[]{ll}\|z^{k+1}-z^{*}\|&=\left\|\displaystyle\frac{1}{k+2}[z^{0}-z^{*}]+\displaystyle\frac{k+1}{k+2}[P_{c_{k}}(z^{k})-z^{*}]+\displaystyle\frac{k+1}{k+2}\eta^{k}\right\|\\[8.0pt] &\leq\displaystyle\frac{1}{k+2}\left\|z^{0}-z^{*}\right\|+\displaystyle\frac{k+1}{k+2}\left\|P_{c_{k}}(z^{k})-z^{*}\right\|+\displaystyle\frac{k+1}{k+2}\left\|\eta^{k}\right\|\\[8.0pt] &\leq\displaystyle\frac{1}{k+2}\left\|z^{0}-z^{*}\right\|+\displaystyle\frac{k+1}{k+2}\left\|z^{k}-z^{*}\right\|+\displaystyle\frac{k+1}{k+2}\left\|\eta^{k}\right\|\\[8.0pt] &\leq\displaystyle\frac{1}{k+2}\left\|z^{0}-z^{*}\right\|+\displaystyle\frac{k+1}{k+2}\left(\|z^{0}-z^{*}\|+\displaystyle\sum_{j=0}^{k-1}\varepsilon_{j}\right)+\displaystyle\frac{k+1}{k+2}\varepsilon_{k}\\[8.0pt] &\leq\|z^{0}-z^{*}\|+\displaystyle\sum_{j=0}^{k}\varepsilon_{j}.\end{array}$
This means that the result holds for $k+1$. The proof is completed. ∎
Proposition 2.1.
Let $\{z^{k}\}$ be any sequence generated by the Halpern accelerated proximal point algorithm under criterion (A) with $\{c_{k}\}$ bounded away from zero. Then $\{z^{k}\}$ is bounded if and only if $T^{-1}(0)\neq\emptyset$.
Proof.
If $T^{-1}(0)\neq\emptyset$, then the boundedness of $\{z^{k}\}$ follows from Lemma 2.2. For the rest of the proof, we assume that $\{z^{k}\}$ is any bounded sequence satisfying (A). Let $s>0$ be such that
$\|z^{k}\|<s\mbox{ and }\varepsilon_{k}<s\quad\forall k.$
Then $\{z^{k}\}$ has at least one weak cluster point $z^{\infty}$ and $\|z^{\infty}\|\leq s$. It follows from (2.1) that
$z^{k+1}=\displaystyle\frac{1}{k+2}z^{0}+\displaystyle\frac{k+1}{k+2}\bar{z}^{k},$
which implies
$\|\bar{z}^{k}\|=\left\|\displaystyle\frac{1+2}{k+1}z^{k+1}-\displaystyle\frac{1}{k+1}z^{0}\right\|\leq 1.4s+\left\|\displaystyle\frac{1}{k+1}z^{0}\right\|<1.5s+0.5s=2s,$
when $k>N_{0}$ for some large integer $N_{0}$. This implies that
$\|P_{c_{k}}(z^{k})\|\leq\|\bar{z}^{k}\|+\varepsilon_{k}\leq 3s,\hbox{ for }k>N_{0}.$
Then, the result follows by applying the same technique as in [10, Theorem 1]. For completeness, the details are provided below. Consider the set-valued mapping $T^{\prime}$ defined by
$T^{\prime}(z)=T(z)+\partial h(z)\quad\forall z\in{\cal H},$
where
${h(z)=\mathbb{I}_{3s\textbf{B}}(z)}\mbox{ where }\textbf{B}=\{z\in{\cal H}:\|z\|\leq 1\}.$
It is easy to obtain
$\partial h(z)=\left\{\begin{array}[]{ll}\{0\},&\mbox{if }\|z\|<3s,\\[4.0pt] \{\lambda z:\lambda\geq 0\},&\mbox{if }\|z\|=3s,\\[4.0pt] \emptyset,&\mbox{if }\|z\|>3s.\end{array}\right$.
Observe that $\partial h$ is a maximal monotone operator, because $h$ is a lower semicontinuous proper convex function; its effective domain is
${\rm Dom}\,\partial h=3s\textbf{B}.$
Furthermore,
$T^{\prime}(z)=T(z)\mbox{ when }z\in{\rm int}\,3s\textbf{B}.$} (2.5)
Since $\|P_{c_{k}}(z^{k})\|<3s$, while
$c_{k}^{-1}(z^{k}-P_{c_{k}}(z^{k}))\in T(P_{c_{k}}(z^{k})).$
So we have
$P_{c_{k}}(z^{k})\in{\rm Dom}\,(T)\cap{\rm int}\,{\rm Dom}\,(\partial h),\quad\forall k>N_{0}.$ (2.6)
$P_{c_{k}}(z^{k})\in(I+c_{k}T^{\prime})^{-1}(z^{k}),\quad\forall k>N_{0}.$ (2.7)
Inasmuch as ${\rm Dom}\,(T)\cap{\rm int}\,{\rm Dom}\,(\partial h)\neq\emptyset$ by (2.6), we know that $T^{\prime}$, as the sum of the maximal monotone operators $T$ and $\partial h$, is itself maximal monotone. Hence $P^{\prime}_{c_{k}}=(I+c_{k}T^{\prime})^{-1}$ is actually single-valued, and then (2.7) implies
$P_{c_{k}}(z^{k})=P^{\prime}_{c_{k}}(z^{k})\quad\mbox{for all large }k>N_{0}.$
Thus the sequence $\{z^{k}\}$ can be regarded as arising from the Halpern accelerated proximal point algorithm for operator $T^{\prime}$. As the effective domian ${\rm Dom}\,(T^{\prime})$ is bounded, so that $(T^{\prime})^{-1}(0)\neq\emptyset$ by Lemma 2.1. It follows from (2.5) that $(T^{\prime})^{-1}(0)\subseteq T^{-1}(0)$, this implies $T^{-1}(0)\neq\emptyset$. The proof is completed. ∎
Theorem 2.1.
Assume that Assumption 2.1 holds. Let $\{z^{k}\}$ and $\{\bar{z}^{k}\}$ be sequences generated by the Halpern accelerated proximal point algorithm under criterion (A) with $c_{k}$ bounded away from zero. Then $\{z^{k}\}$ converges in the weak topology to a point $z^{\infty}$ satisfying $0\in T(z^{\infty})$ and
$\lim_{k\rightarrow\infty}\|Q_{c_{k}}(z^{k})\|=0,\quad\lim_{k\rightarrow\infty}\|\bar{z}^{k}-z^{k}\|=0.$ (2.8)
Proof.
It follows from Proposition 2.1 that the sequence $\{z^{k}\}$ is a bounded sequence because Assumption 2.1 holds. Let $s>0$ be such that
$\|z^{k}\|<s\mbox{ and }\varepsilon_{k}<s,\quad\forall k.$ (2.9)
Just like the proof of Proposition 2.1, there exists an integer $N_{0}>0$ such that
$\|\hat{z}^{k}\|\leq 2s,\quad\|P_{c_{k}}(z^{k})\|\leq 3s,\quad\mbox{ when }k>N_{0}.$
Let $z^{\infty}$ be a weak cluster point of $\{z^{k}\}$ and choose any $\bar{z}\in T^{-1}(0)$. It is easy to obtain the following inequalities
$\|P_{c_{k}}(z^{k})-z^{0}\|^{2}+\|Q_{c_{k}}(z^{k})\|^{2}\leq\|z^{k}-z^{0}\|^{2},\quad\forall k$ (2.10)
and
$\|P_{c_{k}}(z^{k})-\bar{z}\|^{2}+\|Q_{c_{k}}(z^{k})\|^{2}\leq\|z^{k}-\bar{z}\|^{2},\quad\forall k.$ (2.11)
From (2.10), we have
$\begin{array}[]{ll}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-z^{0}\|^{2}+\|z^{k+1}-z^{0}\|^{2}&\leq\|z^{k+1}-z^{0}\|^{2}-\|P_{c_{k}}(z^{k})-z^{0}\|^{2}\\[8.0pt] &=\left(\displaystyle\frac{k+1}{k+2}\right)^{2}\|z^{0}-\bar{z}^{k}\|^{2}-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})+(\bar{z}^{k}-z^{0})\|^{2}\\[8.0pt] &\leq-\|(P_{c_{k}}(z^{k})-\bar{z}^{k})\|^{2}-2\langle(P_{c_{k}}(z^{k})-\bar{z}^{k}),(\bar{z}^{k}-z^{0})\rangle\\[8.0pt] &\leq 2\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|\|\bar{z}^{k}-z^{0}\|\leq 2(\|z^{0}\|+2s)\varepsilon_{k}.\end{array}$
Hence we obtain
$\sum_{k=0}^{\infty}\|Q_{c_{k}}(z^{k})\|^{2}\leq 2(\|z^{0}\|+2s)\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty.$
This implies that $Q_{c_{k}}(z^{k})\rightarrow 0$.
Noting that
$\|z^{k}-\bar{z}^{k}\|=\|z^{k}-P_{c_{k}}(z^{k})+P_{c_{k}}(z^{k})-\bar{z}^{k}\|\leq\|Q_{c_{k}}(z^{k})\|+\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|,$
one has that $\|z^{k}-\bar{z}^{k}\|\rightarrow 0$. From (2.11), we have
$\begin{array}[]{ll}\|Q_{c_{k}}(z^{k})\|^{2}-\|z^{k}-\bar{z}\|^{2}+\|z^{k+1}-\bar{z}\|^{2}&\leq\|z^{k+1}-\bar{z}\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}\|^{2}\\[8.0pt] &=\left\|\displaystyle\frac{1}{k+2}(z^{0}-\bar{z})+\displaystyle\frac{k+1}{k+2}(\bar{z}^{k}-\bar{z})\right\|^{2}-\|P_{c_{k}}(z^{k})-\bar{z}^{k}+\bar{z}^{k}-\bar{z}\|^{2}\\[8.0pt] &\leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}-2\langle P_{c_{k}}(z^{k})-\bar{z}^{k},\bar{z}^{k}-\bar{z}\rangle-\|P_{c_{k}}(z^{k})-\bar{z}^{k}\|^{2}\\[8.0pt] &\leq\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}+2(2s+\|\bar{z}\|)\varepsilon_{k}.\end{array}$
This implies that
$\|z^{k+1}-\bar{z}\|^{2}\leq\|z^{k}-\bar{z}\|^{2}+\displaystyle\frac{1}{(k+2)^{2}}\|(z^{0}-\bar{z})\|^{2}+2(2s+\|\bar{z}\|)\varepsilon_{k},$
which because of
$\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty$
implies the existence of
$\lim_{k\rightarrow\infty}\|z^{k}-\bar{z}\|=\mu<+\infty.$
Since $c_{k}$ is bounded away from zero, one has
$\lim_{k\rightarrow\infty}c_{k}^{-1}\|Q_{c_{k}}(z^{k})\|=0.$ (2.12)
Observe that
$0\leq\langle z-P_{c_{k}}(z^{k}),w-c_{k}^{-1}Q_{c_{k}}(z^{k})\rangle\mbox{ for all }w\in T(z).$ (2.13)
Noting that
$\begin{array}[]{ll}\|z^{k+1}-P_{c_{k}}(z^{k})\|&=\left\|\displaystyle\frac{1}{k+2}z^{0}+\displaystyle\frac{k+1}{k+2}\bar{z}^{k}-P_{c_{k}}(z^{k})\right\|\\[6.0pt] &\leq\left\|\bar{z}^{k}-P_{c_{k}}(z^{k})\right\|+\displaystyle\frac{1}{k+2}\left\|z^{0}-\bar{z}^{k}\right\|\\[6.0pt] &\leq\varepsilon_{k}+\displaystyle\frac{1}{k+2}[2s+\|z^{0}\|],\end{array}$
we have $\|z^{k+1}-P_{c_{k}}(z^{k})\|\rightarrow 0$. Since $z^{\infty}$ is a weak cluster point of $\{z^{k}\}$, it is also a weak cluster point of $\{P_{c_{k}}(z^{k})\}$. Then we get from (2.12) and (2.13) that
$0\leq\langle z-z^{\infty},w\rangle,\quad\forall(z,w)\in{\rm gph}\,T.$
Arguing as in the proof of [10, Theorem 1], one can show that the sequence $\{z^{k}\}$ admits at most one weak cluster point. Hence, the entire sequence $\{z^{k}\}$ converges weakly to some $z^{\infty}$. ∎
3 The rate of convergence
Different from the commonly used technique of performance estimation problems in [5, 2], in this section we adopt a direct algebraic proof to establish the convergence rate of the Halpern-accelerated proximal point algorithm with a fixed parameter $c_{k}\equiv c$. Furthermore, under the regularity condition considered in [10], we show that the accelerated version admits a fast linear convergence rate when the sequence $\{c_{k}\}$ is nondecreasing.
Proposition 3.1.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (A) with $c_{k}\equiv c>0$. Then for $k=1,2,\ldots$
$\|z^{k}-P_{c}(z^{k})\|\leq\sqrt{\displaystyle\frac{4\|z^{0}-z^{*}\|^{2}}{(k+1)^{2}}+\Delta_{k}},$ (3.1)
where
$\begin{array}[]{ll}\Delta_{k}&=\displaystyle\frac{4k}{(k+1)^{2}}\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4k}{k+1}\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$ (3.2)
Proof.
For $\eta^{k}=\bar{z}^{k}-P_{c}(z^{k})$, $1\leq j\leq k$, $z^{j}$ is expressed as
$z^{j}=\displaystyle\frac{1}{j+1}z^{0}+\displaystyle\frac{j}{j+1}(P_{c}(z^{j-1})+\eta^{j-1})\mbox{ or }P_{c}(z^{j-1})=\displaystyle\frac{j+1}{j}z^{j}-\displaystyle\frac{1}{j}z^{0}-\eta^{j-1}.$ (3.3)
By nonexpansiveness of $P_{c}$, we have
$\|P_{c}(z^{k})-z^{*}\|^{2}\leq\|z^{k}-z^{*}\|^{2}\quad\mbox{ for }z^{*}\in T^{-1}(0)$ (3.4)
and
$\|P_{c}(z^{j})-P_{c}(z^{j-1})\|^{2}\leq\|z^{j}-z^{j-1}\|^{2}\quad\mbox{ for }j=1,\ldots,k.$ (3.5)
Making a weighted sum of (3.5), we obtain
$0\geq\displaystyle\sum_{j=1}^{k}j(j+1)\left(\|P_{c}(z^{j})-P_{c}(z^{j-1})\|^{2}-\|z^{j}-z^{j-1}\|^{2}\right).$ (3.6)
In view of the second expression of (3.3), we have
$\begin{array}[]{l}j(j+1)\|P_{c}(z^{j})-P_{c}(z^{j-1})\|^{2}\\[4.0pt] =j(j+1)\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]-\eta^{j-1}\right\|^{2}\\[4.0pt] =j(j+1)\left[\left\|z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}]\right\|^{2}+\|\eta^{j-1}\|^{2}-2\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle\right]\\[4.0pt] =j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2}+2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle+\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}\\[8.0pt] \quad-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2}\\[6.0pt] =A_{1}(j)+A_{2}(j)+A_{3}(j)-2j(j+1)\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\rangle+j(j+1)\|\eta^{j-1}\|^{2},\end{array}$ (3.7)
where
$\begin{array}[]{l}A_{1}(j)=j(j+1)\|z^{j}-P_{c}(z^{j})\|^{2},\,\,A_{2}(j)=2(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-z^{0}\rangle,\,\,\hbox{and}\,\,A_{3}(j)=\displaystyle\frac{j+1}{j}\|z^{j}-z^{0}\|^{2}.\end{array}$
In view of the first expression of (3.3), we have
$\begin{array}[]{l}-j(j+1)\|z^{j}-z^{j-1}\|^{2}\\[4.0pt] =-j(j+1)\left\|\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})+\displaystyle\frac{j}{j+1}\eta^{j-1}-z^{j-1}\right\|^{2}\\[8.0pt] =-2j\langle z^{0}-P_{c}(z^{j-1}),P_{c}(z^{j-1})-z^{j-1}\rangle-j(j+1)\|P_{c}(z^{j-1})-z^{j-1}\|^{2}\\[8.0pt] \quad-\displaystyle\frac{j}{j+1}\|z^{0}-P_{c}(z^{j-1})\|^{2}-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[8.0pt] \quad-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] =B_{1}(j)+B_{2}(j)+B_{3}(j)-j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] \quad-2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle,\end{array}$ (3.8)
where
$\begin{array}[]{l}B_{1}(j)=-\displaystyle\frac{j}{j+1}\|z^{0}-P_{c}(z^{j-1})\|^{2},\,B_{2}(j)=-2j\langle z^{0}-P_{c}(z^{j-1}),P_{c}(z^{j-1})-z^{j-1}\rangle,\\[6.0pt] \hbox{and}\,\,B_{3}(j)=-j(j+1)\|P_{c}(z^{j-1})-z^{j-1}\|^{2}.\end{array}$
We may write $B_{1}(j)$ as
$\begin{array}[]{ll}B_{1}(j)&=-\displaystyle\frac{j}{j+1}\|z^{0}-P_{c}(z^{j-1})\|^{2}\\[6.0pt] &=-\displaystyle\frac{j}{j+1}\left\|\displaystyle\frac{j+1}{j}z^{0}-\displaystyle\frac{j+1}{j}z^{j}+\eta^{j-1}\right\|^{2}\\[6.0pt] &=-\displaystyle\frac{j+1}{j}\|z^{0}-z^{j}\|^{2}-2\langle z^{0}-z^{j},\eta^{j-1}\rangle-\displaystyle\frac{j}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=-A_{3}(j)-2\langle z^{0}-z^{j},\eta^{j-1}\rangle-\displaystyle\frac{j}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$ (3.9)
And we make a sum of $B_{2}(j)$:
$\sum_{j=1}^{k}B_{2}(j)=-\displaystyle\sum_{j=1}^{k}2j\langle z^{0}-P_{c}(z^{j-1}),P_{c}(z^{j-1})-z^{j-1}\rangle=\displaystyle\sum_{j=0}^{k-1}2(j+1)\langle z^{0}-P_{c}(z^{j}),z^{j}-P_{c}(z^{j})\rangle.$
Thus summing up $A_{2}(j)$ and $B_{2}(j)$ for $j=1,\ldots,k$ results in
$\begin{array}[]{l}\displaystyle\sum_{j=1}^{k}[A_{2}(j)+B_{2}(j)]=2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle\\[6.0pt] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\quad+2\displaystyle\sum_{j=1}^{k-1}(j+1)\langle z^{j}-P_{c}(z^{j}),z^{j}-P_{c}(z^{j})\rangle+2\|z^{0}-P_{c}(z^{0})\|^{2}.\end{array}$ (3.10)
Shifting the index in the summation of $B_{3}(j)$
$\sum_{j=1}^{k}B_{3}(j)=-\displaystyle\sum_{j=1}^{k}j(j+1)\|P_{c}(z^{j-1})-z^{j-1}\|^{2}=-\sum_{j=0}^{k-1}(j+1)(j+2)\|P_{c}(z^{j})-z^{j}\|^{2}$
and summing up $A_{1}(j)$ and $B_{3}(j)$ for $j=1,\ldots,k$ gives
$\sum_{j=1}^{k}[A_{1}(j)+B_{3}(j)]=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}-2\displaystyle\sum_{j=1}^{k-1}(j+1)\|z^{j}-P_{c}(z^{j})\|^{2}-2\|z^{0}-P_{c}(z^{0})\|^{2}.$ (3.11)
Thus, inserting (3.9), (3.10) and (3.11) in (3.6) leads to
$\begin{array}[]{ll}0\geq&k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle\\[6.0pt] &-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\right\rangle+\displaystyle\sum_{j=1}^{k}j(j+1)\|\eta^{j-1}\|^{2}\\[6.0pt] &-2\displaystyle\sum_{j=1}^{k}\langle z^{0}-z^{j},\eta^{j-1}\rangle-\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j}{j+1}\|\eta^{j-1}\|^{2}-\displaystyle\sum_{j=1}^{k}j(j+1)\left(\displaystyle\frac{j}{j+1}\right)^{2}\|\eta^{j-1}\|^{2}\\[8.0pt] &-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[12.0pt] =&k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle\\[5.0pt] &+C_{1}+C_{2}+C_{3}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2},\end{array}$ (3.12)
where
$\begin{array}[]{l}C_{1}=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\right\rangle,\\[6.0pt] C_{2}=-2\displaystyle\sum_{j=1}^{k}\langle z^{0}-z^{j},\eta^{j-1}\rangle,\\[6.0pt] C_{3}=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle.\end{array}$
We express $C_{1}+C_{2}$ as
$\begin{array}[]{ll}C_{1}+C_{2}&=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle z^{j}-P_{c}(z^{j})+\displaystyle\frac{1}{j}[z^{j}-z^{0}],\eta^{j-1}\right\rangle+C_{2}\\[6.0pt] &=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle z^{j}-P_{c}(z^{j}),\eta^{j-1}\right\rangle-\displaystyle\sum_{j=1}^{k}2(j+1)\left\langle z^{j}-z^{0},\eta^{j-1}\right\rangle+C_{2}\\[6.0pt] &=D_{1}+D_{2},\end{array}$ (3.13)
where
$\begin{array}[]{l}D_{1}=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle z^{j}-P_{c}(z^{j}),\eta^{j-1}\right\rangle\,\,\hbox{and}\,\,D_{2}=-\displaystyle\sum_{j=1}^{k}2j\left\langle z^{j}-z^{0},\eta^{j-1}\right\rangle.\end{array}$
We further express $C_{3}$ as
$\begin{array}[]{ll}C_{3}&=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle\displaystyle\frac{1}{j+1}(z^{0}-P_{c}(z^{j-1}))+P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[6.0pt] &=-\displaystyle\sum_{j=1}^{k}2j(j+1)\left\langle\displaystyle\frac{1}{j+1}z^{0}+\displaystyle\frac{j}{j+1}P_{c}(z^{j-1})-\displaystyle\frac{j}{j+1}z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[6.0pt] &=E_{1}+E_{2},\end{array}$ (3.14)
where
$\begin{array}[]{l}E_{1}=-2\displaystyle\sum_{j=1}^{k}j\left\langle z^{0}-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\,\,\hbox{and}\,\,E_{2}=-2\displaystyle\sum_{j=1}^{k}j^{2}\left\langle P_{c}(z^{j-1})-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle.\end{array}$
For notational simplicity, we set $\eta^{-1}=0$, then $E_{1}$ can be expressed as
$\begin{array}[]{ll}E_{1}&=-2\displaystyle\sum_{j=1}^{k}j\left\langle z^{0}-z^{j-1},\displaystyle\frac{j}{j+1}\eta^{j-1}\right\rangle\\[6.0pt] &=-2\displaystyle\sum_{j=0}^{k-1}(j+1)\left\langle z^{0}-z^{j},\displaystyle\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &=-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\eta^{j-1}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle.\end{array}$
Thus $D_{2}+E_{1}$ is expressed
$D_{2}+E_{1}=-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle.$ (3.15)
From the definition of $E_{2}$, we have
[…]
Summing up $D_{1}$ and $E_{2}$, we obtain
$\begin{array}[]{ll}D_{1}+E_{2}&=-2k(k+1)\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle+2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad-2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle.\end{array}$ (3.16)
Combining (3.12) with (3.13)–(3.16), we have
$\begin{array}[]{rl}0&\geq k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(C_{1}+C_{2})+C_{3}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+D_{1}+D_{2}+E_{1}+E_{2}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(D_{1}+E_{2})+(D_{2}+E_{1})+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle\\[6.0pt] &\quad-2k(k+1)\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle+2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad-2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$
Dividing the above inequality by $k+1$ and then adding (3.4), we obtain
$\begin{array}[]{ll}0\geq&\displaystyle\frac{k+1}{2}\|z^{k}-P_{c}(z^{k})\|^{2}-\displaystyle\frac{2}{k+1}\|z^{0}-z^{*}\|^{2}+\displaystyle\frac{2}{k+1}\left\|z^{0}-z^{*}-\displaystyle\frac{k+1}{2}(z^{k}-P_{c}(z^{k}))\right\|^{2}\\[6.0pt] &-2\displaystyle\frac{k}{k+1}\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle-\displaystyle\frac{2}{k+1}\displaystyle\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &-\displaystyle\frac{2}{k+1}\displaystyle\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle-2k\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &+\displaystyle\frac{2}{k+1}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &-\displaystyle\frac{2}{k+1}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle+\displaystyle\frac{2}{k+1}\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$
This implies
$\begin{array}[]{ll}\|z^{k}-P_{c}(z^{k})\|^{2}&\leq\displaystyle\frac{4}{(k+1)^{2}}\|z^{0}-z^{*}\|^{2}+\displaystyle\frac{4k}{(k+1)^{2}}\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\\[6.0pt] &+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle+\displaystyle\frac{4k}{k+1}\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle+\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\\[6.0pt] &-\displaystyle\frac{4}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k}\displaystyle\frac{j^{2}}{j+1}\|\eta^{j-1}\|^{2},\end{array}$
which completes the proof. ∎
We next derive an explicit upper bound for $\Delta_{k}$ in (3.1). Let
$\beta_{0}=\sum_{j=0}^{\infty}\varepsilon_{j},\quad\kappa_{0}=2\|z^{0}-z^{*}\|+2\beta_{0}.$
From (2.4), it follows that
$\|z^{k}-z^{0}\|\leq\|z^{k}-z^{*}\|+\|z^{0}-z^{*}\|\leq 2\|z^{0}-z^{*}\|+\beta_{0}<\kappa_{0},$ (3.17)
and
$\|P_{c_{k}}(z^{k})-z^{k}\|\leq 2\|z^{k}-z^{*}\|\leq 2\|z^{0}-z^{*}\|+2\beta_{0}=\kappa_{0}.$ (3.18)
These inequalities provide uniform bounds that will facilitate the estimation of $\Delta_{k}$ in (3.1).
Proposition 3.2.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (A) with $c_{k}\equiv c>0$. Then
$\begin{array}[]{ll}\Delta_{k}&\leq 4\kappa_{0}\varepsilon_{k-1}\displaystyle\frac{k}{(k+1)^{2}}+8\kappa_{0}\beta_{0}\displaystyle\frac{1}{(k+1)^{2}}\\[8.0pt] &\quad+8\kappa_{0}\displaystyle\frac{1}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k-1}j\varepsilon_{j-1}+4\kappa_{0}\varepsilon_{k-1}\\[8.0pt] &\quad+8\kappa_{0}\displaystyle\frac{1}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k-1}\displaystyle(j+1)^{2}\varepsilon_{j-1}.\end{array}$ (3.19)
Proof.
Using (3.17) and (3.18), we can derive estimates for each term in (3.1), which together yield (3.19). For notational simplicity, we set $\varepsilon_{-1}:=\varepsilon_{0}$. Specifically, the estimates can be summarized as follows:
  • (1)
    $\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle\leq\displaystyle\|z^{k}-z^{0}\|\cdot\|\eta^{k-1}\|\leq\kappa_{0}\varepsilon_{k-1}$;
  • (2)
    $\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\leq\displaystyle\displaystyle\sum_{j=0}^{k-1}\|z^{0}-z^{j}\|\cdot\|\frac{j+1}{j+2}\eta^{j}\|\leq\kappa_{0}\beta_{0};$
  • (3)
    $\sum_{j=0}^{k-1}j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle\leq\displaystyle\sum_{j=0}^{k-1}j\|z^{0}-z^{j}\|[\|\eta^{j}\|+\|\eta^{j-1}\|]\leq 2\kappa_{0}\displaystyle\sum_{j=0}^{k-1}j\varepsilon_{j-1};$
  • (4)
    $\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\leq\|z^{k}-P_{c}(z^{k})\|\cdot\|\eta^{k-1}\|\leq\displaystyle\kappa_{0}\varepsilon_{k-1};$
  • (5)
    $\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\leq\|P_{c}(z^{j})-z^{j}\|[\|\eta^{j-1}\|+\|\eta^{j}\|]\leq 2\kappa_{0}\varepsilon_{j-1};$
  • (6)
    $\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\leq\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\|P_{c}(z^{j})-z^{j}\|\|\eta^{j-1}\|\leq\displaystyle\kappa_{0}\displaystyle\sum_{j=0}^{k-1}\displaystyle\varepsilon_{j-1}\leq\displaystyle\kappa_{0}\beta_{0}.$
These estimates, together with the expression of $\Delta_{k}$ in (3.2), imply the desired conclusion. ∎
To analyze the convergence rate of the proposed method under inexact computations, we next specify the tolerance sequence $\{\varepsilon_{k}\}$ in criterion (A). A proper choice of $\{\varepsilon_{k}\}$ not only ensures the summability condition required by Theorem 2.1, but also allows us to obtain a quantitative convergence rate estimate. For criterion (A), we choose
$\varepsilon_{k}=\displaystyle\frac{1}{(k+2)^{1+\delta}},\quad\delta>0,k=0,1,\ldots.$ (3.20)
Then
$\beta_{0}=\beta_{0}(\delta)=\displaystyle\sum_{k=0}^{\infty}\frac{1}{(k+2)^{1+\delta}},\quad\delta>0$ (3.21)
and
$\kappa_{0}=\kappa_{0}(\delta)=2(\beta_{0}(\delta)+\|z^{0}-z^{*}\|),\quad\delta>0.$ (3.22)
Next, we show that with the specific choice of the tolerance, the Halpern’s acceleration proximal point method can achieve the convergence quite close to the tight convergence rate in [6].
Theorem 3.1.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (3.20) with $c_{k}\equiv c>0$. Then for $k=1,2,\ldots$
$\|z^{k}-P_{c}(z^{k})\|\leq\displaystyle\frac{2\|z^{0}-z^{*}\|}{(k+1)}+\sqrt{\Theta_{k}},$ (3.23)
where
[…]
where $\beta_{0}(\delta)$ and $\kappa_{0}(\delta)$ are defined by (3.21) and (3.22), respectively.
Proof.
By substituting the specific choice of $\varepsilon_{k}$ in (3.20) into the definition of $\Delta_{k}$ in (3.19), we have
$\begin{array}[]{cl}\Delta_{k}&\leq 4\kappa_{0}\displaystyle\frac{1}{(k+1)^{2+\delta}}+8\kappa_{0}\beta_{0}\displaystyle\frac{1}{(k+1)^{2}}+A+4\kappa_{0}\displaystyle\frac{1}{(k+1)^{1+\delta}},\end{array}$ (3.24)
where,
$A:=8\kappa_{0}\displaystyle\frac{1}{(k+1)^{2}}\displaystyle\sum_{j=1}^{k-1}\left(\displaystyle\frac{1}{(j+1)^{\delta}}+\displaystyle\frac{1}{(j+1)^{\delta-1}}\right)<8\kappa_{0}\displaystyle\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt.$
To proceed, we estimate the integrals in the above expression by considering different cases for the parameter $\delta>0$. Specifically,
  • (1)
    For $0<\delta<1$, we have
    $\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt<\displaystyle\frac{1}{(1-\delta)(k+1)^{1+\delta}}+\displaystyle\frac{1}{(2-\delta)(k+1)^{\delta}};$
  • (2)
    For $\delta=1$, we have
    $\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt=\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-1}+1)dt<\displaystyle\frac{\ln(k+1)}{(k+1)^{2}}+\displaystyle\frac{1}{(k+1)};$
  • (3)
    For $1<\delta<2$, we have
    $\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt<\displaystyle\frac{1}{(\delta-1)(k+1)^{2}}+\displaystyle\frac{1}{(2-\delta)(k+1)^{\delta}};$
  • (4)
    For $\delta=2$, we have
    $\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt<\displaystyle\frac{1}{(k+1)^{2}}+\frac{\ln(k+1)}{(k+1)^{2}};$
  • (5)
    For $\delta>2$, we have
    $\frac{1}{(k+1)^{2}}\int_{1}^{k+1}(t^{-\delta}+t^{1-\delta})dt<\displaystyle\frac{1}{(\delta-1)(k+1)^{2}}+\frac{\ln(k+1)}{(\delta-2)(k+1)^{2}}.$
These estimates, together with (3.24), complete the proof. ∎
Corollary 3.1.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (3.20) with $c_{k}\equiv c>0$. Then for $k\in\textbf{N}_{+}$,
$\|z^{k}-P_{c}(z^{k})\|\leq\left\{\begin{array}[]{ll}{\rm O}\left(\sqrt{\displaystyle\frac{8\kappa_{0}}{2-\delta}}\displaystyle\frac{1}{(k+1)^{\delta/2}}\right),&\mbox{if }0<\delta<2,\\[12.0pt] {\rm O}\left(\sqrt{8\kappa_{0}\displaystyle\frac{\ln(k+1)}{(k+1)^{2}}}\right),&\mbox{if }\delta=2,\\[12.0pt] {\rm O}\left(\sqrt{8\kappa_{0}\left(\displaystyle\frac{1}{\delta-1}+\frac{1}{\delta-2}+\beta_{0}\right)+4\|z^{0}-z^{*}\|^{2}}\displaystyle\frac{1}{k+1}\right),&\mbox{if }\delta>2.\end{array}\right$.
Remark 3.1.
Corollary 3.1 provides explicit convergence rate estimates for the Halpern accelerated proximal point method under the tolerance rule (3.20). It shows that the residual $\|z^{k}-P_{c}(z^{k})\|$, which measures the violation of the monotone inclusion, converges to zero at a sublinear rate depending on the decay parameter $\delta$. Specifically, the rate improves as $\delta$ increases: for $0<\delta<2$, the decay follows ${\cal O}(k^{-\delta/2})$; at the critical case $\delta=2$, a logarithmic factor appears; and for $\delta>2$, the rate stabilizes at the optimal order ${\cal O}(1/k)$. These results demonstrate that, with appropriately chosen tolerances, the accelerated inexact PPA can approach the theoretically tight convergence behavior established in [6].
We shall say that $T^{-1}$ is Lipschitz continuous at $0\in{\cal H}$ with modulus $a\geq 0$ if there exists a unique solution $\bar{z}$ to $0\in T(z)$ and for some $\bar{\varepsilon}>0$ we have
$\|z-\bar{z}\|\leq a\|w\|\mbox{ whenever }z\in T^{-1}(w)\mbox{ and }\|w\|\leq\bar{\varepsilon}.$ (3.25)
Under the same regularity conditions as those used by Rockafellar [9], as discussed above, we now show that the accelerated version of the proximal point method still attains a fast linear convergence rate. Based on the results in Theorem 2.1, we can derive the following results in a manner similar to that of [10, Theorem 2], and thus, we provide only a sketch of the proof.
Theorem 3.2.
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (B) of (2.2) with $c_{k}$ nondecreasing ($c_{k}\nearrow c_{\infty}\leq+\infty$). Assume that $\{z^{k}\}$ is bounded and $T^{-1}$ is Lipschitz continuous at 0 with modulus $a$; let
$\mu_{k}=\displaystyle\frac{a}{(a^{2}+c_{k}^{2})^{1/2}}<1.$
Then $\{z^{k}\}$ converges strongly to $\bar{z}$, the unique solution to $0\in T(z)$. Moreover, there is an index $\bar{k}$ such that
$\|\bar{z}^{k}-\bar{z}\|\leq\vartheta_{k}\|z^{k}-\bar{z}\|\quad\mbox{ for all }k\geq\bar{k},$ (3.26)
where
$1>\vartheta_{k}=\displaystyle\frac{\mu_{k}+\delta_{k}}{1-\delta_{k}}\geq 0\quad\mbox{ for all }k\geq\bar{k},$
$\vartheta_{k}\rightarrow\mu_{\infty}\quad(\mbox{where }\mu_{\infty}=0\mbox{ if }c_{\infty}=+\infty).$
Proof.
The sequence $\{z^{k}\}$, being bounded, also satisfies criterion (A) of (2.2) for $\varepsilon_{k}=\delta_{k}\|\bar{z}^{k}-z^{k}\|$. Thus the conclusions of Theorem 2.1 hold. Using the same technique in [10, Theorem 2], we can obtain that, there exists a sufficient large number $\bar{k}$ such that
$\|P_{c_{k}}(z^{k})-\bar{z}\|\leq\mu_{k}\|z^{k}-\bar{z}\|\quad\mbox{ if }k\geq\bar{k}.$ (3.27)
Obviously we have
$\|\bar{z}^{k}-\bar{z}\|\leq\|\bar{z}^{k}-P_{c_{k}}(z^{k})\|+\|P_{c_{k}}(z^{k})-\bar{z}\|,$
where under criterion (B) of (2.2)
$\|\bar{z}^{k}-P_{c_{k}}(z^{k})\|\leq\delta_{k}\|\bar{z}^{k}-z^{k}\|\leq\delta_{k}\|\bar{z}^{k}-\bar{z}\|+\delta_{k}\|z^{k}-\bar{z}\|.$
Therefore by (3.27),
$\|\bar{z}^{k}-\bar{z}\|\leq\delta_{k}\|\bar{z}^{k}-\bar{z}\|+\delta_{k}\|z^{k}-\bar{z}\|+\mu_{k}\|z^{k}-\bar{z}\|.$
This inequality produces the one in (3.26) if $\bar{k}\geq\tilde{k}$ is chosen so that (3.26) holds, as is possible since $1>\mu_{k}\searrow 0$ and $\delta_{k}\rightarrow 0$.
Next we have from
$z^{k+1}=\displaystyle\frac{1}{k+2}z^{0}+\displaystyle\frac{k+1}{k+2}\bar{z}^{k}$
that
$\|z^{k+1}-\bar{z}\|\leq\displaystyle\frac{1}{k+2}\|z^{0}-\bar{z}\|+\displaystyle\frac{k+1}{k+2}\|\bar{z}^{k}-\bar{z}\|\rightarrow 0,$
when $k\rightarrow\infty$, this implies the strong convergence of $\{z^{k}\}$ to $\bar{z}$. The proof is completed. ∎
4 Accelerated inexact augmented Lagrange method
Building upon the accelerated inexact proximal point method developed in the previous section, we now extend our analysis to its application in constrained convex optimization within the augmented Lagrangian framework. In analogy to Rockafellar’s seminal work [10, 9], the proposed accelerated inexact proximal point method naturally gives rise to an accelerated inexact augmented Lagrangian method. This connection enables us to establish convergence rate and iteration complexity results for the accelerated iALM in a manner parallel to the analysis of the classical inexact PPA. Throughout this section, we follow the notations in [9].
Consider the optimization problem
$(P)\quad\quad\quad\begin{array}[]{ll}\displaystyle\min_{x\in C}&f_{0}(x)\\[4.0pt] {\rm s.t.}&f_{j}(x)\leq 0,j=1,\ldots,m,\end{array}$ (4.1)
where $C\subseteq\mathbb{R}^{n}$ is a nonempty closed convex set, $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a lower semicontinuous convex function for $i=0,1,\ldots,m$. The Lagrange dual problem of Problem (4.1) is of the form
$(D)\quad\quad\quad\begin{array}[]{ll}\max&g_{0}(y)\\[6.0pt] {\rm s.t.}&y\geq 0,\end{array}$
where
$g_{0}(y)=\inf_{x\in C}\left\{f_{0}(x)+\displaystyle\sum_{j=1}^{m}y_{j}f_{j}(x)\right\}.$
Define the ordinary Lagrangian function $l$ for Problem (4.1) as
$l(x,y)=\left\{\begin{array}[]{ll}f_{0}(x)+\displaystyle\sum_{j=1}^{m}y_{j}f_{j}(x)&\mbox{if }x\in C\mbox{ and }y\in\mathbb{R}^{m}_{+},\\[6.0pt] -\infty&\mbox{if }x\in C\mbox{ and }y\notin\mathbb{R}^{m}_{+},\\[6.0pt] +\infty&\mbox{if }x\notin C.\end{array}\right$.
The essential objective function in the Lagrange dual problem (D) is
$g(y)=\inf_{x\in\mathbb{R}^{n}}l(x,y)=\left\{\begin{array}[]{ll}g_{0}(y)&\mbox{if }y\in\mathbb{R}^{m}_{+},\\[6.0pt] -\infty&\mbox{otherwise}.\end{array}\right$.
It follows from [9] that the augmented Lagrange method for solving (4.1) is the proximal point algorithm to the inclusion
$0\in\partial g(y).$ (4.2)
The proximal mapping associated with $\partial g$, denoted by $P_{c}g$, is defined by
$(P_{c}g)(y)=[I+c\partial g]^{-1}(y),$ (4.3)
which is the unique solution to
$\max_{u}\,g(u)-\displaystyle\frac{1}{2c}\|u-y\|^{2}.$ (4.4)
Now we analyze how to obtain $(P_{c}g)(y)$. For convenience of expression, we use $\tilde{u}=(P_{c}g)(y)$, then $\tilde{u}$ is the unique solution to Problem (4.4). Then for
$Y(x,y,c)=\Pi_{\mathbb{R}^{m}_{+}}[y+cF(x)],$ (4.5)
we have
$\begin{array}[]{ll}g(\tilde{u})-\displaystyle\frac{1}{2c}\|\tilde{u}-y\|^{2}&=\displaystyle\max_{u}\,\left\{g(u)-\displaystyle\frac{1}{2c}\|u-y\|^{2}\right\}\\[6.0pt] &=\displaystyle\max_{u}\inf_{x\in\mathbb{R}^{n}}\left\{l(x,u)-\displaystyle\frac{1}{2c}\|u-y\|^{2}\right\}\\[6.0pt] &=\displaystyle\inf_{x\in\mathbb{R}^{n}}\displaystyle\max_{u}\left\{l(x,u)-\displaystyle\frac{1}{2c}\|u-y\|^{2}\right\}\\[6.0pt] &=\displaystyle\inf_{x\in\mathbb{R}^{n}}\left\{l(x,Y(x,y,c))-\displaystyle\frac{1}{2c}\|Y(x,y,c)-y\|^{2}\right\}\\[6.0pt] &=\displaystyle\inf_{x\in\mathbb{R}^{n}}L(x,y,c),\end{array}$ (4.6)
where $L$ is the augmented Lagrange function
$L(x,y,c)=f_{0}(x)+\displaystyle\frac{1}{2c}\left[\|\Pi_{\mathbb{R}^{m}_{+}}[y+cF(x)]\|^{2}-\|y\|^{2}\right].$
Let $x_{c}(y)$ denote the solution to
$\inf_{x\in\mathbb{R}^{n}}L(x,y,c).$ (4.7)
Then
$(P_{c}g)(y)=\Pi_{\mathbb{R}^{m}_{+}}[y+cF(x_{c}(y))].$ (4.8)
In accelerated inexact proximal point algorithm, one needs to find
$\bar{y}^{k}\approx[P_{c_{k}}g](y^{k}).$
From (4.8), we may hope that this approximate solution $\bar{y}^{k}$ can be implemented by solving the problem
$\inf_{x\in\mathbb{R}^{n}}L(x,y^{k},c_{k}).$ (4.9)
approximately.
This idea was originally developed by Rockafellar [9]. Corresponding to criterion (A) in the inexact proximal point framework, Rockafellar proposed the following stopping criterion:
$\begin{array}[]{l}(C)\quad\quad\quad\quad\quad\quad\phi_{k}(x^{k})-\inf\phi_{k}\leq\varepsilon_{k}^{2}/2c_{k},\quad\displaystyle\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty,\end{array}$ (4.10)
where
$\phi_{k}(x)=L(x,y^{k},c_{k}).$ (4.11)
Now we are in a position to state the accelerated augmented Lagrange method:
[…]
As a direct consequence of [9, Proposition 6], we obtain the following result, which reveals the connection between criterion (C) for the inexact ALM and criterion (A) for the inexact PPM.
Corollary 4.1.
For $P_{c}g$ as in (4.3), $\phi_{k}$ as in (4.11), $x^{k+1}$ is generated by
$x^{k+1}\approx\arg\min\phi_{k}(x)$
according to criterion (C) and $\bar{y}^{k}=Y(x^{k+1},y^{k},c_{k})$, one has
$\|\bar{y}^{k}-[P_{c_{k}}g](y^{k})\|\leq\varepsilon_{k},$
namely criterion (A) is satisfied for $\partial g$.
We recall the notions of an asymptotically minimizing sequence and the asymptotic infimum of Problem (4.1), as introduced by Rockafellar [9]. A sequence $\{x^{k}\}\subseteq C$ is said to be asymptotically minimizing for Problem (4.1) if
$\limsup_{k\to+\infty}f_{i}(x^{k})\leq 0,\quad i=1,\ldots,m,$ (4.13)
and if $\limsup_{k\to+\infty}f_{0}(x^{k})$ attains the smallest possible value among all sequences in $C$ satisfying (4.13). This smallest value is referred to as the asymptotic infimum of Problem (4.1), and is denoted by $\operatorname{asym}\inf(P)$.
Theorem 4.1.
Suppose $\sup(D)>$-$\infty$, and let the accelerated augmented Lagrange method (4.12) be executed. If the generated sequence $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ is bounded, then $y^{k}\rightarrow y^{\infty}$, where $y^{\infty}$ is some optimal solution to Problem (D), and $\{x^{k}\}$ is asymptotically minimizing for Problem (P) with
$f_{i}(x^{k+1})\leq\displaystyle\frac{1}{c_{k}}\left\{y^{k+1}_{i}-y^{k}_{i}+\displaystyle\frac{1}{k+1}[y^{k+1}_{i}-y^{0}_{i}]\right\}\rightarrow 0,\quad\mbox{ for }i=1,\ldots,m,$ (4.14)
$f_{0}(x^{k+1})-{\rm asym}\inf(P)\leq\displaystyle\frac{1}{2c_{k}}\left[\varepsilon_{k}^{2}+\|y^{k}\|^{2}-\|y^{k+1}\|^{2}-\displaystyle\frac{2k+3}{(k+1)^{2}}\left\langle y^{k+1}-y^{0},y^{k+1}-\displaystyle\frac{1}{2k+3}y^{0}\right\rangle\right].$ (4.15)
The boundedness of $\{y^{k}\}$ under (C) is actually equivalent to the existence of an optimal solution to Problem (D). It holds if Problem (P) satisfies the Slater condition, in this case one has $\max\,(D)=\inf\,(P)={\rm asym}\,\inf(P)$.
If $\{y^{k}\}$ is bounded and there exists an $\alpha_{0}$ such that the set
$\{x\in\mathbb{R}^{n}:x\mbox{ is feasible for }(P)\mbox{ and }f_{0}(x)\leq\alpha_{0}\}$
is nonempty and bounded, then the sequence $\{x^{k}\}$ is also bounded, and all of its cluster points are optimal solutions to Problem (P).
Proof.
Corollary 4.1 shows that criterion (C) implies criterion (A) for $T=-\partial g$. Thus the accelerated augmented Lagrange method is reduced to the Halpern accelerated proximal point algorithm under criterion (A) to $0\in\partial-g(y)$. Then from Theorem 2.1, we have that the sequence $\{y^{k}\}$ converges to a solution $y^{\infty}$ to $0\in-\partial g(y)$, namely an optimal solution to Problem (D). From the definition of $y^{k+1}$, we have
$\bar{y}^{k}=\displaystyle\frac{k+2}{k+1}y^{k+1}-\displaystyle\frac{1}{k+1}y^{0}$ (4.16)
and thus
$\bar{y}^{k}_{i}-y^{k}_{i}=y^{k+1}_{i}-y^{k}_{i}+\displaystyle\frac{1}{k+1}[y^{k+1}_{i}-y^{0}_{i}]$ (4.17)
and
$\bar{y}^{k}_{i}-y^{k}_{i}=Y_{i}(x^{k+1},y^{k},c_{k})=\max\left\{-y^{k}_{i},c_{k}f_{i}(x^{k+1})\right\}\geq c_{k}f_{i}(x^{k+1})$
and $y^{k+1}-y^{k}\rightarrow 0$ as well as the boundedness of $\{y^{k}\}$, we obtain (4.14) from (4.17).
Now we prove (4.15). Observing that
$\phi_{k}(x)=f_{0}(x)+\displaystyle\frac{1}{2c_{k}}[\|Y(x,y^{k},c_{k}\|^{2}-\|y^{k}\|^{2}]\quad\mbox{ for }x\in C,$
we have
$\phi_{k}(x^{k+1})-f_{0}(x^{k+1})=\displaystyle\frac{1}{2c_{k}}[\|\bar{y}^{k}\|^{2}-\|y^{k}\|^{2}].$ (4.18)
From the definition of $P_{c_{k}}g](y^{k})$, we have
$\begin{array}[]{ll}\displaystyle\inf_{x}\phi_{k}(x)&=\inf_{x}L(x,y^{k},c_{k})\\[6.0pt] &=\max_{y}\left\{g(y)-\displaystyle\frac{1}{2c_{k}}\|y-y^{k}\|^{2}\right\}\\[6.0pt] &=g([P_{c_{k}}g](y^{k}))-\displaystyle\frac{1}{2c_{k}}\|[P_{c_{k}}g](y^{k})-y^{k}\|^{2},\end{array}$
which implies
$\inf\phi_{k}\leq g([P_{c_{k}}g](y^{k}))\leq\max\,(D).$ (4.19)
Combining (4.18) and (4.19) we get
$\begin{array}[]{ll}f_{0}(x^{k+1})-\max(D)&\leq\phi_{k}(x^{k+1})-\inf\phi_{k}+\displaystyle\frac{1}{2c_{k}}[\|y^{k}\|^{2}-\|\bar{y}^{k}\|^{2}]\\[6.0pt] &\leq\displaystyle\frac{1}{2c_{k}}[\varepsilon_{k}^{2}+\|y^{k}\|^{2}-\|\bar{y}^{k}\|^{2}].\end{array}$ (4.20)
In view of (4.16), we have
$\begin{array}[]{ll}\|\bar{y}^{k}\|^{2}&=\left\|y^{k+1}+\displaystyle\frac{1}{k+1}y^{k+1}-\displaystyle\frac{1}{k+1}y^{0}\right\|^{2}\\[7.0pt] &=\|y^{k+1}\|^{2}+\displaystyle\frac{2k+3}{(k+1)^{2}}\left\langle y^{k+1}-y^{0},y^{k+1}-\displaystyle\frac{1}{2k+3}y^{0}\right\rangle.\end{array}$ (4.21)
But every $x\in C$ satisfies
$f_{0}(x)+\displaystyle\sum_{i=1}^{\infty}y^{\infty}_{i}f_{i}(x)\geq\inf_{x}l(x,y^{\infty})=g(y^{\infty})=\max\,(D),$
so that $\max\,(D)\leq{\rm asym}\,\inf\,(P)$. Therefore (4.20) and (4.21) imply (4.15). ∎
Now we discuss the complexity of the accelerated augmented Lagrange method. Define the average point
$\tilde{x}^{k}=\displaystyle\sum_{j=1}^{k}c_{j}x^{j},\quad\mbox{ where }c_{j}=c_{j-1}\left(\displaystyle\sum_{l=1}^{k}c_{l-1}\right)^{-1}.$ (4.22)
Proposition 4.1.
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrange method (4.12) be executed. Then the generated sequences $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ satisfy
$f_{i}(\tilde{x}^{k})\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}c_{j}\right)^{-1}\left\{y^{k}_{i}-y^{0}_{i}+\displaystyle\sum_{j=1}^{k}\frac{2\Delta_{0}}{j}\right\},\quad\mbox{ for }i=1,\ldots,m,$ (4.23)
$f_{0}(\tilde{x}^{k})-{\rm asym}\inf(P)\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}2c_{j}\right)^{-1}\left[\displaystyle\sum_{j=0}^{k-1}\varepsilon_{j}^{2}+\|y^{0}\|^{2}-\|y^{k}\|^{2}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{12\Delta_{0}^{2}}{j}\right],$ (4.24)
where $\Delta_{0}$ is a constant satisfying
$\|y^{k}\|\leq\Delta_{0}\quad\mbox{ for }k=0,1,2,\ldots.$ (4.25)
Proof.
First of all, in view of Theorem 4.1, as the solution set to Problem (D) is nonempty, the sequence $\{y^{k}\}$ is bounded and there exists a positive constant $\Delta_{0}$ satisfying (4.25). From (4.14) in Theorem 4.1, we have
$c_{k-1}f_{i}(x^{k})\leq\left\{y^{k}_{i}-y^{k-1}_{i}+\displaystyle\frac{1}{k}2\Delta_{0}\right\}\quad\mbox{ for }i=1,\ldots,m,$ (4.26)
which implies that
$\sum_{j=1}^{k}c_{j-1}\times\displaystyle\sum_{j=1}^{k}c_{j}f_{i}(x^{j})\leq\left\{y^{k}_{i}-y^{0}_{i}+\displaystyle\sum_{j=1}^{k}\frac{1}{j}2\Delta_{0}\right\}\quad\mbox{ for }i=1,\ldots,m.$
From Jensen’s inequality, we obtain (4.23) from the above inequality.
In view of (4.15), $(2k+3)/(k+1)^{2}\leq 3/(k+1)$ and
$\left\langle y^{k+1}-y^{0},y^{k+1}-\displaystyle\frac{1}{2k+3}y^{0}\right\rangle\leq 4\Delta_{0}^{2},$
we obtain
$c_{j-1}[f_{0}(x^{j})-{\rm asym}\inf(P)]\leq\displaystyle\frac{1}{2}\left[\varepsilon_{j-1}^{2}+\|y^{j-1}\|^{2}-\|y^{j}\|^{2}-\displaystyle\frac{3}{j}\times 4\Delta_{0}^{2}\right],$ (4.27)
which implies
$\sum_{j=1}^{k}c_{j-1}\times\left[\displaystyle\sum_{j=1}^{k}c_{j}f_{0}(x^{j})-{\rm asym}\inf(P)\right]\leq\displaystyle\frac{1}{2}\left[\displaystyle\sum_{j=1}^{k}\varepsilon_{j-1}^{2}+\|y^{0}\|^{2}-\|y^{k}\|^{2}-\displaystyle\sum_{j=1}^{k}\frac{12}{j}\Delta_{0}^{2}\right].$
Again from Jensen’s inequality, we obtain (4.24) from the above inequality. ∎
Based on Proposition 4.1, by considering different choices of the sequence $\{c_{k}\}$, we can further establish the convergence rate and iteration complexity of the accelerated inexact augmented Lagrangian method. The corresponding results are summarized in the following theorem.
Theorem 4.2.
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrange method (4.12) be executed in which $\varepsilon_{k}$ is defined by (3.20), and $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ are generated sequences.
  • (1)
    If $c_{k}$ is a sequence of increasing positive parameters with $c_{0}>0$, then
    $f_{i}(\tilde{x}^{k})\leq\displaystyle\frac{1}{c_{0}k}[3\Delta_{0}+2\ln k]\quad\mbox{ for }i=1,\ldots,m,$ (4.28)
    $f_{0}(\tilde{x}^{k})-{\rm asym}\inf(P)\leq\displaystyle\frac{1}{2c_{0}k}\left[\displaystyle\frac{2(1+\delta)}{1+2\delta}+13\Delta_{0}^{2}+12\Delta_{0}^{2}\ln k\right]$ (4.29)
  • (2)
    If $c_{k}=c_{0}(k+1)$ with $c_{0}>0$, then
    $f_{i}(\tilde{x}^{k})\leq\displaystyle\frac{2}{c_{0}k(k+1)}[3\Delta_{0}+2\ln k]\quad\mbox{ for }i=1,\ldots,m,$ (4.30)
    $f_{0}(\widetilde{x}^{k})-{\rm asym}\inf(P)\leq\displaystyle\frac{1}{c_{0}k(k+1)}\left[\displaystyle\frac{2(1+\delta)}{1+2\delta}+13\Delta_{0}^{2}+12\Delta_{0}^{2}\ln k\right]$ (4.31)
where $\Delta_{0}$ is a constant (4.25).
Proof.
From (4.23), we have for $i=1,\ldots,m$ that
$f_{i}(\tilde{x}^{k})\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}c_{j}\right)^{-1}\left\{3\Delta_{0}+\displaystyle\sum_{j=2}^{k}\frac{2\Delta_{0}}{j}\right\}\leq\left(kc_{0}\right)^{-1}\left\{3\Delta_{0}+\displaystyle\int_{1}^{k}\frac{2\Delta_{0}}{t}dt\right\}\leq\displaystyle\frac{1}{c_{0}k}[3\Delta_{0}+2\ln k],$
which is just (4.28). From (4.24) and (3.20), we have that
$\begin{array}[]{ll}f_{0}(\tilde{x}^{k})-{\rm asym}\inf(P)&\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}2c_{j}\right)^{-1}\left[\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{1}{(j+1)^{2+2\delta}}+13\Delta_{0}^{2}+\displaystyle\sum_{j=2}^{k}\displaystyle\frac{12\Delta_{0}^{2}}{j}\right]\\[16.0pt] &\leq\displaystyle\left(2c_{0}k\right)^{-1}\left[\displaystyle\sum_{j=1}^{k-1}\displaystyle\frac{1}{(j+1)^{2+2\delta}}+1+13\Delta_{0}^{2}+\displaystyle\sum_{j=2}^{k}\displaystyle\frac{12\Delta_{0}^{2}}{j}\right]\\[16.0pt] &\leq\displaystyle\left(2c_{0}k\right)^{-1}\left[\displaystyle\int_{1}^{k}\displaystyle\frac{1}{t^{2+2\delta}}dt+1+13\Delta_{0}^{2}+\displaystyle\int_{1}^{k}\displaystyle\frac{12\Delta_{0}^{2}}{t}dt\right]\\[16.0pt] &\leq\displaystyle\left(2c_{0}k\right)^{-1}\left[\displaystyle\frac{1}{1+2\delta}+1+13\Delta_{0}^{2}+\displaystyle 12\Delta_{0}^{2}\ln k\right]\\[16.0pt] &=\left(2c_{0}k\right)^{-1}\left[\displaystyle\frac{2(1+\delta)}{1+2\delta}+13\Delta_{0}^{2}+12\Delta_{0}^{2}\ln k\right].\end{array}$
This proves (4.29). Inequalities (4.30) and (4.31) are easily obtained in a similar way. ∎